OpenStudy (anonymous):

The perimeter of a rectangular garden is 344 feet. If the length of the garden is 95 feet, what is its width?

OpenStudy (anonymous):

\[P=2l+2w\]\[w=\frac{p-2l}{2}\]\[w=\frac{244-2(95)}{2}\]\[P=27feet\]

OpenStudy (texaschic101):

P = 2L + 2W P = 344 L = 95 now sub 344 = 2(95) + 2W 344 = 190 + 2W 344 - 190 = 2W 154 = 2W 154/2 = W 77 = W check.. P = 2L + 2W 344 = 2(95) + 2(77) 344 = 190 + 154 344 = 344 (correct) width is 77 ft

OpenStudy (texaschic101):

you could also set it up like twvogels did, except he subbed the wrong number in for perimeter.